OpenStudy (anonymous):

If ∠PLA and ∠ELA are complementary, ∠PLA = 5x – 2, and ∠ELA = x + 8, what is the measure of ∠ELA?

OpenStudy (raden):

when both angles add up to 90 degrees then solve for x ∠PLA + ∠ELA = 90 5x - 2 + x + 8 = 90 6x + 6 = 90 6x = 90 - 6 6x = 84 x = 84/6 x = 14 therefore, the measure of ∠ELA = x+8 = 14+8 = 22 degrees